You have an order to infuse Heparin at 1500 units/hour


Question 1 You have an order to infuse Heparin at 1500 units/hour. The pharmacy has supplied a 250 mL IV bag of D5W with 25,000 units of heparin. Calculate the mL/hr to set the pump 


Question 2 The physician orders lorazepam (Ativan) 1.5 mg L.M. The lorazepam is supplied in a vial with a concentration of 2 mg/mL. How many ml should the nurse give? 


Question 3 2 pts A nurse drew up 3.75 cc's of lorazepam (Ativan). The lorazepam is supplied in a vial containing 2 mg/mL How many mg does the nurse have in the syringe?
Answer #1

Q.1.

250 ml IV bag contains 25000 units of Heparin

order of heparin infusion=1500units/hour

25000 in 250 ml

1 in 250/25000

1500 in=(250/25000) X 1500= 15ml

pump setting ml/hr= 15ml /hr

Q. 2.

Physician order of lorazepam= 1.5 mg IM

supplied 2mg/ml

2mg in 1 ml

1 mg=1/2 ml

1.5 mg in= 1.5/2=0.75 ml

The ml required for ordered value=0.75ml

Q.3. nurse drew up 3.75 cc means ml lorazepam

supplied 2mg/ml

1 ml contains 2mg

3.75 ml will contain=3.75X 2= 7.5 mg

how many mg the nurse have in the syringe by drewing the 3.75ml= 7.5 mg